OpenStudy (anonymous):

Properties of Equaliity, How do I start this off? 6 = -10 - p/5

ganeshie8 (ganeshie8):

u trying to solve it for p value ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

so.. we need to do something and isolate p

OpenStudy (anonymous):

I wanna say multiplication...?

ganeshie8 (ganeshie8):

wow.. u r right lets multiply both sides with 5

OpenStudy (anonymous):

Okay I got 30 = -50 - p/5? Is that right?

ganeshie8 (ganeshie8):

\[5 * 6 = 5(-10) - 5*\frac{p}{5}\]

ganeshie8 (ganeshie8):

\[30 = -50 - p \]

ganeshie8 (ganeshie8):

our goal is to isolate p.. we need to get rid of 50... what can we do now ?

OpenStudy (anonymous):

Add like terms?

ganeshie8 (ganeshie8):

yep... by adding which "-50" gets cancelled ?

ganeshie8 (ganeshie8):

i meant, by adding which number..

OpenStudy (anonymous):

-50 + 50 is now a 0 so I add -50 to 30?

ganeshie8 (ganeshie8):

any number we add, we need to add to both sides. thats the "property of equality" we cant just add to LHS, and equate it to RHS

ganeshie8 (ganeshie8):

think again... which number will cancel out "-50" ?

ganeshie8 (ganeshie8):

30 = -50 - p

OpenStudy (anonymous):

It's not the 30 but the 5?

ganeshie8 (ganeshie8):

if we add "-50" to both sides, RHS becomes "-100 - p" !!!

ganeshie8 (ganeshie8):

its the "-50" we want to cancel below : 30 = -50 -p

ganeshie8 (ganeshie8):

so.. if we add "50" both sides.... 50 + 30 = 50-50-p

ganeshie8 (ganeshie8):

its cancelling out right.. on RHS... u see ?

OpenStudy (anonymous):

Yup, It's 80 = -p/5 now?

ganeshie8 (ganeshie8):

no.. there is no 5 on bottom of "-p"

ganeshie8 (ganeshie8):

80 = -p multiply "-1" both sides -80 = p p = 80

OpenStudy (anonymous):

Oh okay I get it, thank you for your time!

ganeshie8 (ganeshie8):

is it good for you that explanation... or you still uneasy ?

ganeshie8 (ganeshie8):

yw.. :)